OpenStudy (anonymous):

two thirds g + 6=-12

OpenStudy (anonymous):

2/3g=-18

OpenStudy (anonymous):

how do you get that

OpenStudy (anonymous):

subtract 6 from both sides

OpenStudy (anonymous):

the next step is to divide by 2/3

OpenStudy (anonymous):

what about the two thirds tho

OpenStudy (anonymous):

oh so when i get -18 now i have too divide by two thirds? and thats my final answer?

OpenStudy (anonymous):

yes maam

OpenStudy (anonymous):

-27?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

oh thank you sooo much! i really appreciate it!

OpenStudy (anonymous):

you gettin it?

OpenStudy (anonymous):

ok i have one more question lol

OpenStudy (anonymous):

if its in fraction form like c-5 over 4 = 3 how do you do it that way?

OpenStudy (anonymous):

multiply both sides by 4

OpenStudy (anonymous):

then you tell me

OpenStudy (anonymous):

-6?

OpenStudy (anonymous):

how did you get that

OpenStudy (anonymous):

i added -4 too the 4 and 3 then i got -1 and i did -1-5

OpenStudy (anonymous):

you have (c-5)/4=3 then multiply both sides by 4 so c-5 =12

OpenStudy (anonymous):

oh! i get it lol wow. i feel dumb.

OpenStudy (anonymous):

multiplication undoes division, addition undoes subtraction

OpenStudy (anonymous):

oh ok! thank you! you have been a great help!

OpenStudy (anonymous):

youre welcome hope you get it

OpenStudy (anonymous):

i do get it! thanks!

OpenStudy (anonymous):

k